Two forces act on an object. The first is a horizontal 50N force. The second is a 40 N force acting at an angle to the horizontal. Use a graphical method drawn to a scale of 1mm:1N to find the resultant if the 40N force is inclined at 30° to the horizontal



1
Expert's answer
2022-01-30T13:51:23-0500


The sesultant force has a componets

"R_x=85\:N,\quad R_y=20\:N"

The magnitude of resultant force

"R=\sqrt{R_x^2+R_y^2}=\sqrt{85^2+20^2}=87\: N"